OpenStudy (anonymous):

show how to solve 3x^2+10x=8 by factoring.

OpenStudy (alexandervonhumboldt2):

easy: first subtract 8 from both sides: we get 3x^2+10x-8=0 3x^2+12x-2x-8=0 3x(x+4)−2(x+4)=0 (x+4)(3x−2)=0 now find the roots:

OpenStudy (anonymous):

so then wouldn't I make it two separate equations: (x+4)=0 and(3x-2)=0 the solve?

OpenStudy (alexandervonhumboldt2):

yes yo do

OpenStudy (alexandervonhumboldt2):

you equal both parentheses to 0 and solve.

OpenStudy (anonymous):

thank you

OpenStudy (alexandervonhumboldt2):

:) since you are new here, Welcome to OpenStudy!!! You can read this tutorial to learn how to use openstudy. http://openstudy.com/updates/5655f2d1e4b0af360a7cfaa3